OpenStudy (anonymous):

Which of the following is not equivalent to 36/52? 9/13, 18/26, 72/104, 108/154

OpenStudy (anonymous):

\[\LARGE \frac{36\div 2}{52\div 2}=\frac{18\div 2}{26\div 2}=\frac{9}{13}\] so Option A is not. Try option B...

OpenStudy (anonymous):

18/26=9/13 option B is not